# Masked Language Modeling Becomes Conditional Density Estimation for Tabular Data Synthesis

### v1

@@ 2.2 Masked Conditional Density Estimation (MaCoDE)
1:for $j=randperm\{1,2,\cdots,p\}$ do
2:     $\hat{{\bf y}}_{j}\sim Cat\big{(}\pi_{j}(\hat{{\bf y}}\odot\hat{{\bf m}};\theta)/\tau\big{)}$
3:     $\hat{{\bf m}}_{j}\leftarrow 1$
4:for $j=1,2,\cdots,p$ do
5:     if $j\in I_{C}$ then
6:         $\hat{{\bf x}}_{j}\leftarrow\hat{F}^{-1}_{j}(u)$, where $u\sim U(b_{\hat{{\bf y}}_{j}-1},b_{\hat{{\bf y}}_{j}})$
7:     if $j\in I_{D}$ then
8:         $\hat{{\bf x}}_{j}\leftarrow\hat{{\bf y}}_{j}$
2.3 Theoretical Results
We consider a factorization of the ground-truth joint PDF $p^{*}({\bf x}_{1},\cdots,{\bf x}_{p})$ according to an arbitrary permutation $\sigma=[\sigma_{1},\sigma_{2},\cdots,\sigma_{p}]$ of the indices $\{1,2,\cdots,p\}$:
$p^{*}({\bf x}_{1},\cdots,{\bf x}_{p})$ $=$ $p^{*}_{\sigma_{1}}({\bf x}_{\sigma_{1}})\cdot p^{*}_{\sigma_{2}}({\bf x}_{\sigma_{2}}|{\bf x}_{\sigma_{1}})\cdot p^{*}_{\sigma_{3}}({\bf x}_{\sigma_{3}}|{\bf x}_{\sigma_{1}},{\bf x}_{\sigma_{2}})\cdots p^{*}_{\sigma_{p}}({\bf x}_{\sigma_{p}}|{\bf x}_{\sigma_{1}},\cdots,{\bf x}_{\sigma_{p-1}}).$
As discussed in Remark 1, our objective function (3) facilitates the estimation of $p_{\sigma_{j}}^{*}({\bf x}_{\sigma_{j}}|{\bf x}_{\sigma_{1}},\cdots,{\bf x}_{\sigma_{j-1}})$ for any $\sigma$ and $j$ by utilizing the mask distribution of Definition 2. Without loss of generality, let $\sigma$ be an identity permutation. We define the masking vector ${\bf m}^{(j)}$ as follows:
${\bf m}^{(j)}_{k}$$\coloneqq$ $\begin{cases}1,&\text{if $k\in\{1,\cdots,j-1\}$}\\ 0,&\text{otherwise}\end{cases},$
and for notational simplicity, let ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$ for all $j$. For $j\in I_{C}$, under Assumption 1,
$p_{j}^{*}({\bf x}_{j}|{\bf x}_{-j})$ $=$ $c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot p^{*}_{j}({\bf x}_{j})$ (4)
@@ 2.3 Theoretical Results
Assumption 2 (Smoothness).
For all $j\in I_{C}$, $c^{*}_{j}(\cdot|\cdot)$ is twice differentiable, ${{c_{j}^{*}}^{\prime}}(\cdot|\cdot)$ is continuous, and $({{c_{j}^{*}}^{\prime}}(\cdot|\cdot))^{2}$ is finite.
In this subsection, to establish the asymptotic property, we denote the number of bins as $L_{n}$, indicating its dependency on the sample size $n$. We define the histogram-based density estimator for $c^{*}_{j}$:
$\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)$$\coloneqq$ $\sum_{l=1}^{L_{n}}\frac{\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))}{1/L_{n}}\cdot\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta),$
where $\tilde{F}=\{\tilde{F}_{j}:j\in I_{C}\}$ denotes another set of CDF estimators with density function $\tilde{p}_{j}$. Although only $\hat{F}$ is used in the training process of MaCoDE (see Algorithm 1), we introduce two different sets of CDF estimators, $\hat{F}$ and $\tilde{F}$, to demonstrate the theoretical properties of MaCoDE.
Assumption 3 (Model assumptions).
For all $j\in I_{C}$, ($\pi_{j}(\cdot;\theta)$, $\hat{F}_{j}$) and $\tilde{F}_{j}$ are independent, and $\tilde{p}_{j}$ is kernel density estimator. $L_{n}\rightarrow\infty$ and $L_{n}/n\rightarrow 0$ as $n\rightarrow\infty$. For all $j\in I_{C}$, $\sup_{l\in[L_{n}]}\mbox{Bias}(\pi_{jl})=\mathcal{O}(n^{\alpha_{1}}L_{n}^{\alpha_{2}})$, $\sup_{l\in[L_{n}]}\mbox{Var}(\pi_{jl})=\mathcal{O}(n^{\beta_{1}}L_{n}^{\beta_{2}})$, and $\sup_{l\in[L_{n}]}\mbox{Var}\big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\big{)}=\mathcal{O}(n^{\gamma_{1}}L_{n}^{\gamma_{2}})$ with $\alpha_{1}+\alpha_{2}+1\leq 0$, $\beta_{1}+\beta_{2}+3\leq 0$, and $\gamma_{1}+\gamma_{2}+3\leq 0$.
We satisfy Assumption 3 by partitioning the $n$ training samples into two equal halves, each containing $n/2$ training samples. Subsequently, $\pi_{j}(\cdot;\theta)$ and $\hat{F}$ are trained using the first half of the training samples, while $\tilde{F}_{j}$ is fitted using the second half of the training samples.
Theorem 1 (Consistency).
Under Assumption 1, 2 and 3, for all $j\in I_{C}$,
$\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)$$\coloneqq$ $\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j}),$ (5)
is weakly consistent estimator for $p_{j}^{*}({\bf x}_{j}|{\bf x}_{-j})$.
Synthetic samples can be generated from (5) as follows: Firstly, perform histogram sampling from $\hat{c}_{j}$ by choosing a histogram bin according to $\pi_{jl}$ and uniformly sampling from that bin interval. Then, apply inverse transform sampling of $\tilde{p}_{j}$ with respect to the output of histogram sampling. Therefore, Theorem 1 shows that our estimator $\hat{p}_{j}$ is weakly consistent, which implies that Algorithm 2 can theoretically generate asymptotically valid synthetic samples for continuous columns. Although we do not have findings that the deep learning estimator $\pi_{jl}$ satisfies Assumptions 3, a multi-class linear logistic regression satisfies the conditions specified in Assumption 3 [16, 29].
Remark 2.
Since we constrain the values of continuous columns to the interval $[0,1]$, our proposed approach and Theorem 1 can be seen as similar to the copula density estimation method [22, 27, 9, 2]. However, given the heterogeneous nature of tabular data [3, 43], handling categorical variables is essential, and the categorical marginal distributions do not guarantee the uniqueness of the copula [44]. Therefore, instead of utilizing the copula density estimation method, we adopt a multi-class classification approach for density estimation.
2.4 With Missing Data
@@
Suppose the pattern of missingness varies individually for each observation and is defined by a corresponding missing indicator, denoted as ${\bf r}$, where the indicators for all observations are represented as $\{{\bf r}^{(i)}\}_{i=1}^{n}$. Here, ${\bf r}_{j}=0$ indicates that ${\bf x}_{j}$ is missing, while ${\bf r}_{j}=1$ denotes that ${\bf x}_{j}$ is observed. In cases where the training data contains missing entries (e.g., not a number), it is not feasible to input these missing entries and minimize their log-likelihoods. Therefore, to handle missing value inputs using $g$, irrespective of the value of ${\bf x}_{j}$, we further define $g$ with respect to Definition 1 as follows: for any $F$, $g({\bf x};F)_{j}\coloneqq 0$ if ${\bf r}_{j}=0$. This indicates the bin index ‘0’ is assigned to a missing value input.
@@
${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)\coloneqq\sum_{j:{\bf m}_{j}=0,{\bf r}_{j}=1}\sum_{l=1}^{L}\mathbb{I}(g({\bf x};\hat{F})_{j}=l)\cdot\log\pi_{jl}(g({\bf x};\hat{F})\odot\min({\bf m},{\bf r});\theta),$ (6)
$\min({\bf m},{\bf r})$ is element-wise minimum operation, and $\hat{F}$ is estimated using the observed dataset (refer to [6]). Note that, in equation (6), we do not minimize the log-likelihood of missing entries.
@@
In (6), the missingness pattern is described by $\min({\bf m},{\bf r})$, indicating that the missing data model is determined by the joint distribution $p({\bf m},{\bf r}|{\bf x})$. Thus, as the masking vector defined in Definition 2 satisfies the conditions in Proposition 1, the missingness pattern of our proposed model also conforms to the MAR mechanism according to Proposition 1 if the given data is MAR. It implies that our inference strategies based on the observed dataset can be justified [34]. However, we empirically demonstrate in Section 3 that our proposed model is also applicable to other missing data scenarios.
@@ 3 Experiments
3.1 Overview
Questions. We conduct experiments in which we can provide answers to the following three experimental questions:
@@ 3.1 Overview
Datasets. Similar to several recent studies [18, 34, 37, 36, 54, 20], we utilize 10 real tabular UCI and Kaggle [^1] https://archive.ics.uci.edu/, https://www.kaggle.com/datasets/ datasets of varying sizes. We split the dataset into training and testing sets with a ratio of 80% for training and 20% for testing. Detailed statistics of these datasets are provided in Appendix A.3. Note that we include covtype dataset, which comprises approximately 580K rows, to demonstrate the scalability of our proposed model. We set $L=50$ and $\tau=1$ for all datasets to assess the generalizability of our proposed model unless otherwise stated. [^2] We run experiments using NVIDIA A10 GPU, and our experimental codes are available with pytorch.
Baseline models. For Q1 and Q2, we compare MaCoDE with CTGAN [51], TVAE [51], CTAB-GAN [56], CTAB-GAN+ [57], DistVAE [1], TabDDPM [24], and TabMT [13]. For Q3, we selected the following multiple imputation models that can handle heterogeneous tabular datasets: MICE [48], GAIN [52], missMDA [21], VAEAC [19], MIWAE [34], not-MIWAE [18], and EGC [55]. Detailed experimental settings for these baseline models are provided in Appendix A.5.
Additional evaluations. Due to space limitations, comprehensive experimental settings and results concerning privacy preservability and sensitivity analysis, varying the temperature parameter $\tau$ and missingness rate, are presented in Appendices A.6.1 and A.6.2, respectively.
Table 1: Q1: Statistical fidelity and machine learning utility. The means and the standard errors of the mean across 10 datasets and 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better. The best value is bolded, and the second best is underlined.
@@
3.2 Evaluation Metrics
For all metrics, we report the mean and the standard error of the mean (error bar) across 10 different random seeds and 10 datasets.
@@ 3.2 Evaluation Metrics
Q1. To evaluate the quality of generated synthetic data, we employ two metrics: statistical fidelity [40] and machine learning utility [15]. For statistical fidelity, we utilize the Kullback–Leibler divergence (KL) and the Goodness-of-Fit (GoF) test (continuous: the two-sample Kolmogorov-Smirnov test statistic, categorical: the Chi-Squared test statistic) to assess marginal distributional similarity. Additionally, we employ the Maximum Mean Discrepancy (MMD) and 2-Wasserstein distance (WD) to measure joint distributional similarity. Regarding machine learning utility, we adopt four specific metrics outlined in [15]: regression performance (SMAPE, symmetric mean absolute percentage error), classification performance ($F_{1}$), model selection performance (Model), and feature selection performance (Feature). Refer to Appendix A.5 for a detailed evaluation procedure.
@@
Q2. To ascertain the capability of our proposed model to generate high-quality synthetic datasets even when the given training dataset contains missing values, we evaluate our proposed model, trained using masked training datasets (i.e., incomplete datasets), in downstream regression and classification tasks on the test dataset. We utilize the SMAPE metric for regression tasks and the $F_{1}$ score for classification tasks. See Appendix A.5 for a detailed evaluation procedure.
Q3. We assess the effectiveness of multiple imputations by employing interval inference for the population mean, which was proposed by Rubin [41]. We report the bias, coverage, and confidence interval length, and the evaluation procedure for multiple imputations is outlined in Algorithm 3 [47]. Following [36, 20, 54], we generate the missing value mask for each dataset with three mechanisms (MCAR, MAR, MNAR) in four settings. See Appendix A.5 for detailed missing value generation mechanisms. For all missing data mechanisms, unless otherwise specified, the missingness rate is set to 30%.
Remark 3.
We acknowledge that existing imputation methods have been evaluated using RMSE (root mean square error), a metric for assessing single imputation methods [18, 34, 37, 36, 54, 20]. However, since our objective focuses on distributional learning rather than recovering missing values, we adopt the evaluation procedure proposed in [41, 47], better suited for assessing multiple imputation methods.
3.3 Results

### v2

@@ 3 Proposal
1:  for $j=randperm\{1,2,\cdots,p\}$ do {tabular data lacks the inherent ordering between columns}
2:     $\hat{{\bf y}}_{j}\sim Cat\big{(}\pi_{j}(\hat{{\bf y}}\odot\hat{{\bf m}};\theta)/\tau\big{)}$choosing a histogram bin (index) according to $\pi_{j}$}
3:     $\hat{{\bf m}}_{j}\leftarrow 1$un-mask the $j$th column}
4:  end for
5:  for $j=1,2,\cdots,p$ do
6:     if $j\in I_{C}$ then
7:        $u\sim U(b_{\hat{{\bf y}}_{j}-1},b_{\hat{{\bf y}}_{j}})$uniform sampling a quantile level from the bin interval}
8:        $\hat{{\bf x}}_{j}\leftarrow\hat{F}^{-1}_{j}(u)$inverse transform sampling of $\hat{p}_{j}$}
9:     end if
10:     if $j\in I_{D}$ then
11:        $\hat{{\bf x}}_{j}\leftarrow\hat{{\bf y}}_{j}$
12:     end if
13:  end for
Synthetic data generation. Tabular data lacks the inherent ordering between columns, unlike natural language. Therefore, as outlined in Algorithm 1, MaCoDE randomly generates one column at a time, conditioned on masked subset sizes from $p$ to $1$, in descending order ($p\rightarrow p-1\rightarrow\cdots\rightarrow 2\rightarrow 1$).
Controllable privacy level. Adjusting the privacy level during synthetic data generation is crucial in tabular domain (Park et al. 2018). Similar to TabMT (Gulati and Roysdon 2023), we can also regulate the privacy level using a single hyper-parameter $\tau$ without the need for re-training, while other existing synthesizers have fixed trade-offs between synthetic data quality and privacy level (Xu et al. 2019; Zhao et al. 2023b; Kotelnikov et al. 2023) or require re-training (An and Jeon 2023).
By increasing the temperature parameter $\tau$ in $\hat{{\bf y}}_{j}\sim Cat\big{(}\pi_{j}(\hat{{\bf y}}\odot\hat{{\bf m}};\theta)/\tau\big{)}$ at line 2 of Algorithm 1, we can mitigate the risk of privacy leakage. Figure 4 shows that MaCoDE allows for a trade-off between synthetic data quality (feature selection performance) and privacy preservability (DCR, Distance to Closest Record) as $\tau$ increases (see the Appendix for detailed results).
Figure 2: Trade-off between quality and privacy. Left: feature selection performance. Right: DCR. Error bars represent standard errors. See the Appendix for detailed results.
Theoretical Results
In this section, we aim to provide theoretical insights into MaCoDE’s capabilities in conditional density estimation. Firstly, we consider a factorization of the ground-truth joint PDF $p^{*}({\bf x}_{1},\cdots,{\bf x}_{p})$ according to an arbitrary permutation $\sigma=[\sigma_{1},\sigma_{2},\cdots,\sigma_{p}]$ of the indices $\{1,2,\cdots,p\}$:
$p^{*}_{\sigma_{1}}({\bf x}_{\sigma_{1}})\cdot p^{*}_{\sigma_{2}}({\bf x}_{\sigma_{2}}|{\bf x}_{\sigma_{1}})\cdot p^{*}_{\sigma_{3}}({\bf x}_{\sigma_{3}}|{\bf x}_{\sigma_{1}},{\bf x}_{\sigma_{2}})$
××$\times\cdots\times p^{*}_{\sigma_{p}}({\bf x}_{\sigma_{p}}|{\bf x}_{\sigma_{1}},\cdots,{\bf x}_{\sigma_{p-1}}).$
As discussed in Remark 1, our objective function (5) facilitates the estimation of $p_{\sigma_{j}}^{*}({\bf x}_{\sigma_{j}}|{\bf x}_{\sigma_{1}},\cdots,{\bf x}_{\sigma_{j-1}})$ for any $\sigma$ and $j$ by utilizing the mask distribution of Definition 2.
Without loss of generality, in this section, let $\sigma$ be an identity permutation. Since $c^{*}_{j}$ is the conditional density of $\hat{F}_{j}({\bf x}_{j})$, the conditional probability of $\hat{F}_{j}({\bf x}_{j})$ in the $l$th bin, $\pi_{jl}^{*}({\bf x}_{-j})$, is defined as
@@
where ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$.
Assumption 2.
For all $j$, there exists $K_{j}\geq 0$ such that $c^{*}_{j}$ is $K_{j}$-Lipschitz.
Assumption 2 implies that $c^{*}_{j}$ is constrained in its rate of change, which allows us to estimate the conditional density using a piece-wise constant function (Wasserman 2006; Tsybakov and Tsybakov 2009).
Based on (2), by the change of variable under Assumption 1, we define our conditional density estimator for $p^{*}_{j}$ as follows:
$\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)=\hat{c}_{j}(\hat{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\hat{p}({\bf x}_{j})$ (6)
$=$ $\left(\sum_{l=1}^{L}\frac{\mathbb{I}(\hat{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))}{1/L}\cdot\pi_{jl}\right)\cdot\hat{p}({\bf x}_{j}),$
where $\pi_{jl}$ denotes $\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)$ and ${\bf m}^{(j)}$ is a masking vector such that
${\bf m}^{(j)}_{k}$$\coloneqq$ $\begin{cases}1,&\text{if $k\in\{1,\cdots,j-1\}$}\\ 0,&\text{otherwise}\end{cases}.$
Note that $\hat{c}_{j}$ is the histogram-based conditional density estimator. Synthetic samples can be generated from (6) as follows: Firstly, perform histogram sampling from $\hat{c}_{j}$ by choosing a histogram bin according to $\pi_{jl}$ and uniformly sampling from that bin interval. Then, apply inverse transform sampling of $\hat{p}_{j}$ with respect to the output of histogram sampling. This procedure is outlined in Algorithm 1.
Proposition 1.
Under Assumption 1 and 2,
$\text{TV}\Big{(}p^{*}_{j}(\cdot|{\bf x}_{-j}),\hat{p}_{j}(\cdot|{\bf x}_{-j};\theta)\Big{)}$$\leq$ $\frac{K_{j}}{2L}+\frac{\sqrt{Bias(\theta)}}{\sqrt{2}/L}$
for all $j\in I_{C}$, and ${\bf x}\in\mathbb{R}^{p}$. Here, $\text{TV}(\cdot,\cdot)$ denotes the total variation distance, and $Bias(\theta)$ is defined as:
$Bias(\theta)=\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})$
-- $\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\right]$
where ${\bf y}_{j}|{\bf x}_{-j}$ is a random variable having a categorical distribution such that $\Pr({\bf y}_{j}=l|{\bf x}_{-j})=\Pr(g({\bf x};\hat{F})_{j}=l|{\bf x}_{-j})=\pi_{jl}^{*}({\bf x}_{-j})$ for all $l\in[L]$.
In the definition of $Bias(\theta)$, the second term on the right-hand side corresponds to the classification loss with respect to the target distribution given by (3). This implies that $Bias(\theta)$ can be minimized when the classification loss is minimized. And Proposition 1 demonstrates that the total variation distance between the ground-truth conditional density and our conditional density estimator $\hat{p}_{j}$ is upper bounded by $Bias(\theta)$.
Therefore, minimizing the orderless multi-class classification loss leads to minimizing the total variation distance between conditional distributions, making Algorithm 1 capable of generating theoretically valid synthetic samples under certain assumptions. Note that Proposition 1 holds for any arbitrary permutation $\sigma$, and the CDF transformation is crucial for our proposed method, as it ensures the validity of Proposition 1.
With Missing Data
@@
Suppose the pattern of missingness varies individually for each observation and is defined by a corresponding missing indicator, denoted as ${\bf r}$, where the indicators for all observations are represented as $\{{\bf r}^{(i)}\}_{i=1}^{n}$. Here, ${\bf r}_{j}=0$ indicates that ${\bf x}_{j}$ is missing, while ${\bf r}_{j}=1$ denotes that ${\bf x}_{j}$ is observed. In cases where the training data contains missing entries (e.g., not a number), it is not feasible to input these missing entries and minimize their log-likelihoods. Therefore, to handle missing value inputs using $g$, irrespective of the value of ${\bf x}_{j}$, we further define $g$ for Definition 1 as follows: for any $F$, $g({\bf x};F)_{j}\coloneqq 0$ if ${\bf r}_{j}=0$. This indicates the bin index ‘0’ is assigned to a missing value input.
@@
${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)$$\coloneqq$ $-\sum_{j:{\bf m}_{j}=0,{\bf r}_{j}=1}\sum_{l=1}^{L}\mathbb{I}(g({\bf x};\hat{F})_{j}=l)$
×$\times\log\pi_{jl}(g({\bf x};\hat{F})\odot\min({\bf m},{\bf r});\theta),$
$\min({\bf m},{\bf r})$ is element-wise minimum operation, and $\hat{F}$ is estimated using the observed dataset (Chenouri, Mojirsheibani, and Montazeri 2009). Note that, in ${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)$, we do not minimize the negative log-likelihood of missing entries.
@@
In ${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)$, the missingness pattern is described by $\min({\bf m},{\bf r})$, indicating that the missing data model is determined by the joint distribution $p({\bf m},{\bf r}|{\bf x})$. Thus, as the masking vector defined in Definition 2 satisfies the conditions in Proposition 2, the missingness pattern of our proposed model also conforms to the MAR mechanism according to Proposition 2 if the given data is MAR. It implies that our inference strategies based on the observed dataset can be justified (Mattei and Frellsen 2019). However, we empirically demonstrate in Section 4 that our proposed model is also applicable to other missing data scenarios.
@@ 4 Experiments
Overview
We conduct experiments in which we can provide answers to the following three experimental questions:
@@
MaCoDE(MCAR) - - - - .$168_{\pm.011}.$$623_{\pm.023}.$
MaCoDE(MAR) - - - - .$167_{\pm.011}.$$626_{\pm.023}.$
MaCoDE(MNARL) - - - - .$169_{\pm.011}.$$624_{\pm.023}.$
MaCoDE(MNARQ) - - - - .$164_{\pm.010}.$$630_{\pm.023}.$
Table 1: Q1 and Q2. The means and the standard errors of the mean across 10 datasets and 10 repeated experiments are reported. Across all missingness patterns, a missingness rate of 0.3 is employed. $\uparrow$ ($\downarrow$) denotes higher (lower) is better. The best value is bolded, and the second best is underlined.
Datasets. Similar to several recent studies (Gulati and Roysdon 2023; Kotelnikov et al. 2023), we utilize 10 publicly available real tabular UCI and Kaggle [^1] https://archive.ics.uci.edu/, https://www.kaggle.com/datasets/ datasets of varying sizes and the number of columns. Detailed statistics of these datasets are provided in the Appendix. Note that we include covtype dataset, which comprises approximately 580K rows, to demonstrate the scalability of our proposed model.
Baseline models. For MaCoDE, we set $L=50$ and $\tau=1$ for all datasets. Detailed hyperparameter settings are provided in the Appendix. [^2] We run experiments using NVIDIA A10 GPU, and our experimental codes are available with pytorch. For Q1 and Q2, we compare MaCoDE with CTGAN (Xu et al. 2019), TVAE (Xu et al. 2019), CTAB-GAN (Zhao et al. 2021), CTAB-GAN+ (Zhao et al. 2023b), DistVAE (An and Jeon 2023), TabDDPM (Kotelnikov et al. 2023), and TabMT (Gulati and Roysdon 2023). For Q3, we selected the following multiple imputation models that can handle mixed-type tabular datasets: MICE (van Buuren and Groothuis-Oudshoorn 2011), GAIN (Yoon, Jordon, and van der Schaar 2018), missMDA (Josse and Husson 2016), VAEAC (Ivanov, Figurnov, and Vetrov 2019), MIWAE (Mattei and Frellsen 2019), not-MIWAE (Ipsen, Mattei, and Frellsen 2021), and EGC (Zhao, Townsend, and Udell 2022). Detailed experimental settings for these baseline models are provided in the Appendix.
To assist in interpreting the metrics, we include a baseline synthetic dataset where the synthetic data comprises half of the real training dataset. This dataset, referred to as ‘Baseline,’ serves as a soft upper bound for evaluating the quality of the synthetic data.
Additional evaluations. Due to space constraints, (1) the results for controlling data privacy levels with varying temperature parameter $\tau$, and (2) the results for Q3 (including related works and sensitivity analysis regarding the missingness rate) are provided in the Appendix.
Evaluation Metrics
For all metrics, we report the mean and standard error of the mean (error bars) across 10 different random seeds and 10 datasets. For each random seed, the dataset is randomly split into training and testing sets with an 80% training and 20% testing ratio in the evaluation of Q1 and Q2. During evaluation, the synthetic dataset is generated to have the same number of samples as the real training dataset.
@@
Q1. To evaluate the quality of generated synthetic data, we employ two metrics: statistical fidelity (Qian, Davis, and van der Schaar 2023) and machine learning utility (Hansen et al. 2023). For statistical fidelity, we utilize the Kullback–Leibler divergence (KL) and the Goodness-of-Fit (GoF) test (continuous: the two-sample Kolmogorov-Smirnov test statistic, categorical: the Chi-Squared test statistic) to assess marginal distributional similarity. Additionally, we employ the Maximum Mean Discrepancy (MMD) and 1-Wasserstein distance (WD) to measure joint distributional similarity. These metrics measure how well the synthetic data maintains statistical fidelity to the real training dataset.
Regarding machine learning utility, we use four metrics outlined in (Hansen et al. 2023): regression performance (SMAPE, symmetric mean absolute percentage error), classification performance ($F_{1}$), model selection performance (Model), and feature selection performance (Feature). These metrics are assessed by fitting the machine learning model on the real training and synthetic datasets individually and then comparing the performance of the two models on the test dataset. For a detailed evaluation procedure of Q1, refer to the Appendix.
@@
Q2. To assess our proposed model’s capability to generate high-quality synthetic datasets despite missing values in the training data, we evaluate the model trained on incomplete training datasets. Since other metrics require a complete training dataset for measurement, we only assess synthetic data quality using downstream regression and classification tasks on the test dataset, as these metrics can be evaluated using the test dataset alone, as in Q1. See the Appendix for a detailed evaluation procedure.
Q3.333Since some off-the-shelf packages (missMDA, gcimpute) fail to deliver useful results, we include only 5 datasets that provide meaningful results: abalone, banknote, breast, redwine, and whitewine. We assess the effectiveness of multiple imputations by employing interval inference for the population mean, which was proposed by Rubin (Rubin and Schenker 1986). We report the bias, coverage, and confidence interval length. The detailed evaluation procedure for multiple imputations and the missing value generation mechanisms (MCAR, MAR, MNARL, MNARQ) is provided in the Appendix.
Results